The 5 BS BRDs for installation at LLO are now assembled and tuned in the LLO Optics Lab. The target frequencies are 16.846 Hz and 24.672 Hz (measurements from A. Effler, I will post an alog in the LLO logbook about it).
As per the LHO ones, I tuned them at lower frequencies (-1 or -2%) in order to anticipate the frequency drifts.
For the LHO & LLO bounce modes and for LLO roll modes, I had to put two copper blocks per end at the blade tips in order to reach the target frequency.
There are several factors to take into account:
* Extrapolated from the k CIT experimental k.
I would have expected smaller variations in the spring constant.
The resulting masses are:
Even if the mass kit had the expected weight, I would have not been able to build the BRDs with the masses from the kit (or use several washers of 0.1g).